# A scalable adaptive deep Koopman predictive controller for real-time optimization of mixed traffic flow

### v1

@@ Abstract
The use of connected automated vehicle (CAV) is advocated to mitigate traffic oscillations in mixed traffic flow consisting of CAVs and human driven vehicles (HDVs). This study proposes an adaptive deep Koopman predictive control framework (AdapKoopPC) for regulating mixed traffic flow. Firstly, a Koopman theory-based adaptive trajectory prediction deep network (AdapKoopnet) is designed for modeling HDVs car-following behavior. AdapKoopnet enables the representation of HDVs behavior by a linear model in a high-dimensional space. Secondly, the model predictive control is employed to smooth the mixed traffic flow, where the combination of the linear dynamic model of CAVs and linear prediction blocks from AdapKoopnet is embedded as the predictive model into the AdapKoopPC. Finally, the predictive performance of the prosed AdapKoopnet is verified using the HighD naturalistic driving dataset. Furthermore, the control performance of AdapKoopPC is validated by the numerical simulations. Results demonstrate that the AdapKoopnet provides more accuracy HDVs predicted trajectories than the baseline nonlinear models. Moreover, the proposed AdapKoopPC exhibits more effective control performance with less computation cost compared with baselines in mitigating traffic oscillations, especially at the low CAVs penetration rates. The code of proposed AdapKoopPC is open source111https://github.com/SpaceTrafficSafetyTeam/AdapKoopPC.
Index Terms:
Mixed traffic flow, Koopman operator theory, Predictive control, Data-driven modeling, Real-time optimization.I Introduction
With the acceleration of urbanization and the increase in the number of motor vehicles, traffic congestion and traffic accidents have become increasingly prominent [1]. Against this background, autonomous driving technology has emerged and is expected to become an innovative solution to alleviate traffic congestion[2, 3]. The most common application is Cooperative Adaptive Cruise Control (CACC), which relies on V2V state sharing to enable vehicles within a vehicle platoon to run at coordinated speeds to maintain a small headway [4]. This technique can enhance traffic capacity and improve safety between vehicles, which has attracted a lot of interest from academia and industry. The excellent performance of connected automated vehicles (CAVs) in a purely intelligent network environment has been widely recognized[5]. In essence, CAV is a mobile scanner that provides historical and real-time trajectory data of individual vehicles[6]. As well, it is also a dynamic actuator for adaptive control[7]. In this intelligent network ideal scenario, vehicles communicate seamlessly and autonomous driving systems operate together accurately and efficiently, thus building a highly intelligent transportation ecosystem [8]. However, the industrialization of autonomous driving technology is a complex process, and the realization of large-scale application of mobile internet and autonomous driving technology is a long process. A more realistic traffic environment is that CAVs and human-driven vehicles (HDVs) share the road. This mixed traffic flow prompts scholars to think deeply and study the performance of CAVs in complex traffic environments [9]. Considering the research topics of this paper, we mainly focus on the following two key points, including the modeling of HDVs behavior and the controlling of CAVs.
How to model the HDVs driving behavior in dynamic traffic flow, and to predict the HDVs trajectory with high accuracy? A key goal is to simulate the dynamic driving behaviors of HDVs in mixed traffic flow. These behaviors will be essentially different due to differences in driving environments, drivers, etc. HDV adjusts its own acceleration based on the headway to the leader vehicle, velocity difference, and among others. In term of the physics-driven methods, well-known car-following models were widely employed, such as optimal velocity model [10], full velocity difference model [11], intelligent driver model [12], to describe and model the HDVs behaviors. However, it is known that these time-invariant car-following models always archives high-biased results with large error. With the rapid development of artificial intelligence, big data and computing power, the deep learning methods, including recurrent neural networks (RNN), long short-term memory model (LSTM), deep deterministic policy gradient (DDPG), etc., are being gained more and more interests in modeling the driving behaviors of HDVs. Although deep learning is superior in capturing and modeling driving behavior in complex traffic environments, due to the characteristic of nonlinearity, its shortcomings of slow solution speed and inability to meet real-time optimization control are self-evident.
How to control CAVs more effectively in mixed traffic flows for mitigating traffic oscillations? As well known, the jam-absorption driving strategy in the Newell’s car-following theory framework was proposed to mitigate the propagation of upstream disturbances, and the corresponding vehicles need to be selected based on the expected absorption speed and real-time traffic conditions to implement the strategy [13]. As mentioned before, CAV is a dynamic executor, and its role is no longer limited to“follower”, but has been further expanded to“leader”, which is the basic idea and connotation of the Leading Cruise Control (LCC) framework [14]. In addition, the controllability, observability and head-to-tail stability of the LCC framework were also analyzed and verified, confirming its positive role in improving traffic performance [15, 16]. Along this direction, the Data-Enabled Predictive Leading Cruise Control (Deep-LCC) was proposed [17, 18], as an alternative to parametric car-following model, to achieve safe and optimal control of CAVs in mixed platoons. The Deep-LCC has inspired this paper, however, it still exhibits certain limitations. To be specific, in order to use CAVs to positively guide HDVs, it is necessary to first understand the driving behavior changes of HDV after being affected by CAVs. In Deep-LCC, all HDVs follow a rule-based car-following model and add random high-frequency noise to construct the trajectory database. It is obviously unrealistic and limited to combine these trajectories to match the corresponding HDV expected driving trajectories for different scenarios. this is obviously not realistic and has limitations.
Overall, the dilemmas faced by CAV control for traffic optimization in mixed traffic flow lies in accurately modeling and predicting the driving behavior of HDVs in complex scenarios, as well as meeting the requirements of effective and feasible real-time solution and optimization control in reality. Motivated by the above two requirements, the Koopman operator offers a new solution for identifying and analyzing nonlinear systems. The Koopman operator emerges as a highly suitable approach, which facilitates the handling of intricate nonlinear dynamics through linear transformations. This method allows for an efficient modeling of the HDVs behavior without introducing excessive complexity, and its linear properties are well suited for subsequent real-time optimization of mixed traffic flow. Therefore, this paper develops a multi-scenario adaptive deep Koopman predictive control framework for mixed traffic flow. The main contributions of this paper are as follows:
An adaptive deep Koopman network (AdapKoopnet) is proposed for modeling HDVs car-following behavior with the high-dimensional linear model;
A scalable state prediction model of mixed traffic flow is developed by integrating linear dynamic model of CAVs and linear prediction blocks from AdapKoopnet;
An adaptive deep Koopman predictive control framework (AdapKoopPC) based on the state prediction model is proposed for mitigating traffic oscillations in mixed traffic flow.
The framework of this paper is as follows. Section II mainly reviews the relevant literatures and summarizes the research gap. Section III states the AdapKoopnet. Section IV proposes the AdapKoopPC framework. Section V and Section VI conducts the experiments for HDVs longitudinal trajectory prediction and mixed traffic flow optimization control. At last, the conclusion and prospects are given in Section VII .
II Related work
II-A Driving scenario and characteristics recognition
Recognizing and extracting potential driving scenarios is beneficial for achieving vehicle trajectory prediction with strong generalization ability and accuracy. [19] extracted driving scenarios directly from real driving data and clustered repeated patterns into potential scenario groups without any pre-definition or rules. [20] developed a general active learning framework to annotate driving trajectory time series data and explore unknown driving scenario trajectories. Beyond scenario recognition, more research focuses on capturing driving characteristics. [21] utilized unsupervised algorithms to automatically extract descriptive driving patterns, and the clustered patterns promote a comprehensive understanding of driver behavior characteristics. In addition, identifying the driving style is also a work with practical value. For example, based on the multi-dimensional characteristics, three different driving styles are identified using an unsupervised clustering algorithm[22]. [23] employed a set of experiments to collect driving data among different drivers, and a supervised machine learning-based driving style classifier was designed to recognize the driving style. [24] defined the intermediate variable of neural process as the driving style vector and linked to an interpretable and continuous aggressiveness index. However, identifying and clustering scenarios from natural driving datasets using neural networks without relying on pre-definition or pre-labeling remains to be solved. Secondly, further extracting personalized driving features based on recognized scenarios and using them for driving behavior prediction is also worth exploring.
II-B Car following behavior modeling and trajectory prediction
The trajectory prediction methods of car-following behavior are mainly divided into physics-based methods, data-driven methods, and hybrid-driven methods. The physics-based method uses microscopic traffic flow models for trajectory prediction. For example, [25] explicitly considered the interactions between vehicles and proposed a dynamic Bayesian networks-based filter to estimate the behavior of traffic participants and predict their future trajectories. [26] developed a cooperative intelligent driver model that dynamically determines its acceleration based on parameters such as velocity difference and headway. With the advancement of big data and deep learning technology, data-driven methods for trajectories prediction have been emerging. [27] built spatiotemporal attention long short-term memory (STA-LSTM) for vehicle trajectory prediction, ensuring accuracy while improving interpretability. Pioneeringly, [28] first proposed a physics informed deep learning framework for car-following modeling, taking full advantage of data-driven and physics-based models to surpass existing models. Further, relying on the self-attention mechanism to achieve deeper mining of trajectory features, [29] proposed a Physics-Informed Transformer-Intelligent Driver Model to predict longitudinal vehicle trajectories. However, there is still a contradiction between the complex deep learning model that describes the nonlinear state evolution of vehicle driving behavior and the requirement that CAV control relies on real-time solutions.
II-C CAV control for optimized mixed traffic flow
Various control approaches have been explored to optimize mixed traffic flow performance such as safety, efficiency and energy consumption. [30] presented a highly efficient platoons control framework based on tube MPC to address heavy computational and communication burden, which mitigates prediction uncertainty through feedback control. [31] implemented a cooperative control strategy for CAVs under the assumption of completely random disturbances triggered by HDVs. Recently, a control strategy called “The Follower Stopper controller” was proposed, and real vehicle experiments demonstrated its ability to improve traffic flow stability [32]. To address the limitation of executing FS policies individually for all CAVs, [33] established a cellular automaton model of mixed traffic flow based on the FS strategy for CAV control, which positively affected the efficiency, oscillation, and fuel consumption of mixed traffic flow. From the perspective of collective function optimization, [34] studied the impact of different formations of CAVs on traffic performance, focusing on the formation of cooperation between CAVs using centralized optimal controllers. Besides, a traffic-smoothing controller was directly learned from trajectories based on a reinforcement learning policy gradient algorithm [35]. [36] integrated an advanced linear controller and a DPPO-based DRL controller to construct a hybrid controller, achieving stable and efficient longitudinal driving of CAVs within mixed flow. [37] developed the robust-safety-critical traffic controller in mixed traffic flow. This controller ensures collision-free safety even in the face of actuator and sensor delays, as well as disturbances from the leading HV. As traffic systems expand, a consequent problem is the dramatic increase in online computing burden, and the data-driven distributed control emerged as the times require. Surprisingly, [38] conducted data-driven mixed vehicle platoons dynamics modeling based on Koopman, which is a groundbreaking work that has attracted our high attention and interest. [17, 18] proposed distributed data-driven predictive control that directly utilizes measurable traffic data to design collision-free optimal CAV control inputs for collaboratively smoothing mixed traffic flows. However, it is difficult to demonstrate its generalization ability in other mixed traffic flow scenarios by modeling the entire traffic flow and obtaining the Koopman operator based on a dataset generated by simple physical models, which drives our idea. We expect to use Koopman theory to learn vehicle features from real-world trajectories and construct a scene adaptive and scalable control framework.
II-D The applications of Koopman theory in traffic flow
Koopman operator theory is a mathematical framework that involves representing nonlinear systems in a higher-dimensional linear space, making it possible to study the evolution of complex systems [39] . [40] used Koopman operator to propose a model-free, data-driven approach that cleverly analyzed and predicted the evolution of highly complex nonlinear traffic flow. [41] applied Koopman operator theory and the dynamical mode decomposition for signalized traffic flow networks control, which allows for early identification of unstable queue growth. In order to realize the real-time control of the ramp metering on the freeway efficiently, [42] proposed a model predictive controller with the trained deep Koopman model. The above methods relied on parameter estimation of physical models or modeling of nonlinear dynamics using neural networks. [43] introduced a novel data-driven vehicle modeling and control approach, employing an interpretable Koopman operator base deep neural networks in which extended dynamic mode decomposition was utilized to learn a finite-dimensional approximation of the Koopman operator. Recently, in solving non-stationary time series prediction problems, [44] proposed an efficient time series prediction model driven by Koopman theory, which hierarchically mines the dynamic system of time series.
@@ II-D The applications of Koopman theory in traffic flow
III AdapKoopnet: Adaptive deep Koopman network for car following behavior modeling and prediction of HDVs

### v3

# Mitigating Traffic Oscillations in Mixed Traffic Flow with Scalable Deep Koopman Predictive Control

@@ Abstract
Mitigating traffic oscillations in mixed flows of connected automated vehicles (CAVs) and human-driven vehicles (HDVs) is critical for enhancing traffic stability. A key challenge lies in modeling the nonlinear, heterogeneous behaviors of HDVs within computationally tractable predictive control frameworks. This study proposes an adaptive deep Koopman predictive control framework (AdapKoopPC) to address this issue. The framework features a novel deep Koopman network, AdapKoopnet, which represents complex HDV car-following dynamics as a linear system in a high-dimensional space by adaptively learning from naturalistic data. This learned linear representation is then embedded into a Model Predictive Control (MPC) scheme, enabling real-time, scalable, and optimal control of CAVs. We validate our framework using the HighD dataset and extensive numerical simulations. Results demonstrate that AdapKoopnet achieves superior trajectory prediction accuracy over baseline models. Furthermore, the complete AdapKoopPC controller significantly dampens traffic oscillations with lower computational cost, exhibiting strong performance even at low CAV penetration rates. The proposed framework offers a scalable and data-driven solution for enhancing stability in realistic mixed traffic environments. The code is made publicly available 111https://github.com/SpaceTrafficSafetyTeam/AdapKoopPC.
keywords:
[…]
Traffic oscillations, or stop-and-go waves, are a pervasive phenomenon in congested traffic, leading to significant detriments in traffic efficiency, fuel economy, and road safety (Li et al., 2014; Chen et al., 2014). The advent of Connected and Automated Vehicles (CAVs) offers a transformative approach to mitigate these instabilities (Zhao et al., 2024; Wang et al., 2025). By acting as mobile actuators, CAVs can be strategically controlled to actively dampen the propagation of traffic waves, thereby smoothing the overall traffic flow (Stern et al., 2018; Liu et al., 2025; Zhang et al., 2025b). However, the widespread adoption of CAVs will be a gradual process, leading to a prolonged transition period characterized by mixed traffic, where CAVs must coexist and interact with a majority of human-driven vehicles (HDVs) (Gan et al., 2024; Zou et al., 2025). This complex environment presents a fundamental challenge for effective CAV control (Zhou et al., 2025a; Lai, 2024).
The central challenge lies in a fundamental trade-off between the high-fidelity prediction of HDV behavior and the stringent requirements of real-time, safety-critical vehicle control. On one hand, any proactive control strategy hinges on the ability to accurately anticipate the reactions of surrounding HDVs (Zheng et al., 2024; Jiao et al., 2024). Yet, human driving behavior is inherently nonlinear, stochastic, and heterogeneous, making it difficult to capture with conventional physics-based car-following models (Zhou et al., 2025b; Punzo et al., 2021). While advanced data-driven models, such as those based on deep learning, have demonstrated superior accuracy in capturing these complex dynamics (Hart et al., 2024), their non-convex and computationally intensive nature makes them ill-suited for direct integration into real-time optimization frameworks (Wang et al., 2023a; Long et al., 2024). On the other hand, Model Predictive Control (MPC) stands out as an ideal framework for CAV control, as it systematically handles state and input constraints to guarantee safety and performance (Zhang et al., 2025a; Wang et al., 2023b, 2024c). The efficacy of MPC, however, is predicated on the availability of a computationally tractable—typically linear or convex—predictive model of the entire system, including the HDV responses. This creates a dilemma: high-fidelity nonlinear prediction models compromise real-time control feasibility, whereas overly simplified models fail to capture essential traffic dynamics, leading to suboptimal or unsafe CAV actions (Li et al., 2025; Lyu et al., 2024).
To resolve this prediction-control dilemma, this paper introduces the Koopman operator theory as a powerful mathematical framework to model and predict nonlinear HDV dynamics in a manner that is amenable to real-time predictive control. The Koopman operator provides a global, linear representation of a nonlinear dynamical system by lifting the system’s state into a higher-dimensional observable space (Zhang et al., 2024c; Lusch et al., 2018). This approach allows us to decouple the complexity of modeling from the complexity of control. By leveraging deep neural networks, we can learn the intricate mapping from the original state space to this linear Koopman space, thereby capturing the nonlinear car-following behaviors of HDVs with high fidelity. The resulting model, however, is a linear system in the lifted space, which can be seamlessly and efficiently integrated into an MPC framework. This enables us to achieve both accurate, data-driven prediction and real-time, constrained optimal control.
In this paper, we propose an Adaptive deep Koopman Predictive Control framework (AdapKoopPC) for regulating mixed traffic flow. This framework is designed to learn heterogeneous HDV behaviors from naturalistic driving data and deploy scalable, real-time control strategies for CAVs to mitigate traffic oscillations. The main contributions of this work are threefold:
We develop a deep Koopman network-based trajectory prediction model, termed AdapKoopnet, which adaptively identifies latent driving scenarios and extracts personalized driver characteristics from unlabeled, naturalistic trajectory data to deliver highly accurate HDV behavior predictions.
We propose the scalable AdapKoopPC control framework, which embeds the linear predictive blocks from pre-trained AdapKoopnet models into an MPC formulation. This framework can be dynamically constructed based on the local vehicle composition, enabling real-time, cooperative control of CAVs in arbitrary mixed traffic configurations.
We conduct extensive simulations to validate our framework’s performance. The results confirm that AdapKoopPC markedly reduces traffic oscillations and enhances traffic stability in both small- and large-scale systems (10 and 50 vehicles, respectively), even at low CAV penetration rates, while maintaining computational feasibility for real-time deployment.
The remainder of this paper is organized as follows. Section 2 reviews related literature and highlights the research gap. Section 3 details the architecture of the AdapKoopnet for HDV trajectory prediction. Section 4 presents the AdapKoopPC framework for mixed traffic control. Sections 5 and 6 present experimental validation for the prediction and control models, respectively. Finally, Section 7 concludes the paper and discusses future research directions.
2 Related work
2.1 Car-Following Behavior Modeling and Prediction of HDVs
Accurate prediction of HDVs behavior is a critical prerequisite for the effective control of CAVs in mixed traffic. The challenge is to develop a model that is not only descriptively powerful but also computationally compatible with real-time, optimization-based control frameworks. Historically, HDV modeling has relied on physics-based car-following models, such as the Intelligent Driver Model (IDM) and the Optimal Velocity Model (OVM) (Treiber et al., 2000; Bando et al., 1998; Kim and Yeo, 2024). While these models offer an interpretable structure, their fixed parameters struggle to capture the profound nonlinearity and heterogeneity of real-world driving behavior, often leading to significant prediction errors (Zhang et al., 2024b, a). To address this accuracy deficit, researchers have increasingly adopted data-driven methods. Deep learning models, particularly Long Short-Term Memory (LSTM) networks, excel at learning complex temporal dependencies from trajectory data and have achieved state-of-the-art prediction accuracy (Xu et al., 2024a; Zhang et al., 2019). However, their "black-box" nature and computationally intensive nonlinear structures are fundamentally ill-suited for direct integration into real-time control frameworks, which demand tractable and often linear models for rapid optimization. To further enhance model fidelity, another line of research focuses on personalizing predictions by incorporating human factors. These efforts include identifying distinct driving styles, such as aggressiveness (Zhang et al., 2022; Chen et al., 2024b), or clustering behaviors into different traffic scenarios (Song et al., 2023; Li et al., 2021). While such methods improve predictive power by accounting for driver- or situation-specific nuances, they often introduce additional complexity without resolving the underlying conflict between model nonlinearity and real-time control feasibility. Some hybrid approaches, like Physics-Informed Neural Networks (PINNs), attempt to find a balance by embedding physical laws into the learning process but still typically yield nonlinear representations (Mo et al., 2021; Kang et al., 2023).
Therefore, a critical gap remains in HDV behavior modeling: the need for a framework that can learn complex, personalized driving behaviors from data, yet yields a predictive model with a computationally tractable structure suitable for real-time applications. Existing paradigms force a trade-off between predictive fidelity and computational feasibility, motivating the search for novel modeling approaches that can bridge this divide.
2.2 CAV Control Strategies for Mitigating Traffic Oscillations
Building upon the challenge of accurately modeling HDV behavior, the next critical step is to design effective control strategies for CAVs that can actively mitigate traffic oscillations. The paradigm for such controllers has evolved from localized, reactive maneuvers to holistic, predictive frameworks that optimize the behavior of the entire mixed-traffic system Wang et al. (2023a, c). Early strategies sought to create spatial buffers to absorb traffic waves. These range from pre-planned maneuvers like Jam-Absorption Driving (JAD), where a vehicle is guided to slow down in anticipation of a jam (Nishi et al., 2013; He et al., 2016), to specific control laws like the Follower Stopper (FS) controller, which uses larger-than-normal headways to smooth traffic and has been validated in field experiments (Stern et al., 2018; Wang et al., 2024b). While conceptually insightful, such methods are often limited in their adaptability, motivating a shift towards more proactive and comprehensive control frameworks like Leading Cruise Control (LCC) (Wang et al., 2021). In the LCC framework, a CAV acts as both a follower adapting to preceding vehicles and a leader actively influencing following vehicles, empowering it to exert system-level control to stabilize traffic flow (Zheng et al., 2020; Shang et al., 2024).
The implementation of the LCC concept has been explored through three principal control methodologies, each presenting a distinct trade-off between model fidelity, computational tractability, and safety assurance. (1) MPC is a natural fit for LCC due to its inherent ability to handle complex system dynamics and enforce hard constraints on states and inputs, ensuring safety and comfort (Zhang et al., 2025c). Its performance, however, is fundamentally tethered to the accuracy of the underlying predictive model, which typically requires an explicit, often linearized, model of HDV behavior (Wang et al., 2021). This inevitably creates a mismatch with the true nonlinear dynamics of human drivers, potentially degrading control performance and robustness. (2) Data-Driven Predictive Control aims to circumvent this reliance on explicit models. A prominent example, the Data-EnablEd Predictive Leading Cruise Control (DeeP-LCC) framework, constructs a non-parametric model directly from historical trajectory data (Wang et al., 2023b). While effective, this approach requires large amounts of high-quality offline data that must be representative of online conditions, and the resulting optimization can be computationally demanding for large-scale systems (Wang et al., 2023a; Shang et al., 2024). (3) Reinforcement Learning (RL) offers a powerful model-free alternative, capable of learning complex, nonlinear control policies through interaction with the environment (Li et al., 2024; Xu et al., 2024b; Shi et al., 2025). Despite its adaptability, the principal drawback of RL is the profound difficulty in providing formal safety guarantees. Encouraging safety through reward shaping does not offer assurance against violating hard state constraints, a critical limitation for safety-critical applications like autonomous driving (Zhou et al., 2024).
In summary, the implementation of advanced CAV control faces a methodological trilemma: 1) MPC offers robust constraint handling but is hindered by model mismatch with real-world HDV behavior; 2) Data-driven methods like DeeP-LCC avoid explicit modeling but are data-intensive and computationally demanding; and 3) Reinforcement Learning excels at learning complex policies but struggles to provide formal safety guarantees. This highlights the need for a holistic control framework that synergizes the strengths of these approaches—namely, the data-driven adaptability, the rigorous constraint satisfaction of MPC, and real-time computational efficiency. Our development of a Koopman operator-based predictive control framework is directly motivated by this systemic need.
2.3 The applications of Koopman operator theory in traffic flow
The Koopman operator theory emerges as a powerful framework to reconcile the aforementioned trade-offs by representing nonlinear dynamics with a linear operator in a higher-dimensional space (Wang et al., 2024a; Lusch et al., 2018). Its key advantage lies in constructing an explicit, data-driven linear predictive model from complex trajectory data, thereby bridging the gap between high-fidelity nonlinear models and the computational demands of real-time MPC (Xiao et al., 2022; Chen et al., 2024a). Capitalizing on these advantages, researchers have begun applying Koopman operator theory across various scales of traffic control. At the macroscopic level, it has been used to analyze and forecast complex network-wide traffic dynamics (Avila and Mezić, 2020) and to design efficient controllers for signalized intersections and freeway ramp metering (Ling et al., 2020; Gu et al., 2023; Das et al., 2023). More pertinent to this work, its application has extended to the microscopic level for modeling and controlling vehicle platoons. Studies have successfully employed Koopman-based methods to create data-driven models of mixed-vehicle platoons, combining deep learning with physical interpretability to enhance accuracy and stability analysis (Tian et al., 2024; Zhan et al., 2022). Crucially, recent work has leveraged these learned linear models as the predictive core within MPC frameworks to achieve robust, real-time control of CAVs in mixed traffic (Li et al., 2025; Lyu et al., 2025).
Despite this progress, existing Koopman-based control strategies exhibit two key limitations that hinder their practical scalability. First, their validation often relies on trajectory data generated from idealized car-following simulations, which may not capture the full complexity of real-world driving behaviors. Second, and more critically, current methods typically learn a single, monolithic Koopman operator for a fixed platoon configuration (i.e., a specific number and arrangement of vehicles). This static approach lacks adaptability; if the traffic composition changes, the model becomes invalid. This necessity for a scalable and adaptive framework, capable of learning from naturalistic data and handling dynamic traffic formations, directly motivates the contributions of this paper.
@@ 2.3 The applications of Koopman operator theory in traffic flow
3 AdapKoopnet: Adaptive deep Koopman network for car following behavior modeling and prediction of HDVs